OpenStudy (anonymous):

Algebra 1 9th grade PLEAS EXPLAIN Find the x-intercept of the parabola y=x2+2x-15 using ant representation you prefer. Then explain your method.

OpenStudy (anonymous):

\[y=x ^{2}+2x-15\] This is what the equation looks like

OpenStudy (kinggeorge):

Well, the first step is to set it equal to 0\[0=x^2+2x-15\]Then, you have to factor the equation. Since you have a 1x^2, you;r factorization must look like\[(x+ \text{__})(x+ \text{__})\]where you fill in the blanks with two numbers that multiply to -15 and add to 2. After some quick guess and check, we find that \[5*{-3} = 15, \qquad 5-3 = 2\]so the factorization is\[0=(x+5)(x-3)\]Therefore, one of the following is true if it intercepts the x-axis.\[\begin{matrix} 1. & x+5=0 \\ 2. & x-3=0 \end{matrix}\]Therefore, the x-intercepts are at (-5, 0) and (3, 0). Hope this helps clear up how to do this sort of problem.

OpenStudy (anonymous):

Thanks(: